Session: Oct 2005 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Tao An at Sat Oct 29 16:48:31 2005) Minor failures : RECE
There are something wrong with L-band .The signs of EE and RR have some faults.

Medicina (scalambra at Sat Oct 29 17:37:41 2005) Minor failures : HIGH
v7 and v8 are out of if rec. bandwith

Urumqi (Zhang hua at Sun Oct 30 02:19:07 2005) Minor failures :
Polarization leakage with L-band.start record 10:42(UT),because of power supply trouble.

Torun (Kaz Borkowski at Mon Oct 31 07:10:46 2005) Success :
No known problems.

Westerbork (Tony Foley at Mon Oct 31 08:44:27 2005) Minor failures : RECO RFI PHAS
Wb used in phased array mode. Very bad RFI about 1620MHz Central file server down from 09:03-10:15 UT, so lost the array. Mark5 system froze up, and needed a reboot. so lost 14:51-15:16UT.

Jodrell_Bank (Alastair Gunn at Tue Nov 1 12:25:48 2005) Success :
The Lovell Telescope was used. No known problems.

Effelsberg (A. Kraus at Wed Nov 2 11:43:31 2005) Minor failures : RECO
Data lost between UT1150-1200 due to a problem with the antenna control computer. MK5 problems led to the loss of data between UT1117-1134 and UT1452-1552.

Onsala (Michael Lindqvist at Mon Nov 7 05:44:38 2005) Success :
No known problems.